Assume that thermometer readings are normally distributed with a mean of 0degreesC and a standard deviation of 1.00degreesC. A thermometer is randomly selected and tested. For the case​ below, draw a​ sketch, and find the probability of the reading.​ (The given values are in Celsius​ degrees.)
Between 1.50 and 2.25
The probability of getting a reading between 1.50 degrees C and 2.25 degrees C is?

Answer by rothauserc(4718)

Probability (P) (1.50 < X < 2.25) = P(X < 2.25) -P(X < 1.50)
:
z-score(2.25) = (2.25 -0)/1 = 2.25
:
find in table of z-scores the probability corresponding with a z-score of 2.25
:
P(X < 2.25) = 0.9878
:
z-score(1.50) = (1.50 -0)/1 = 1.50
:
P(X < 1.50) = 0.9332
:
P(1.50 < X < 2.25) = 0.9878 -0.9332 = 0.0546
: